# Towards Principled Multi-Agent Task Agnostic Exploration

### v1

@@ Appendix A Proofs of the Main Theoretical Results
The main empirical proof of concept was based on two environments. First, Env. (i), the so called secret room environment by Liu et al. (2021a). In this environment, two agents operate within two rooms of a $10\times 10$ discrete grid. There is one switch in each room, one in position $(1,9)$ (corner of first room), another in position $(9,1)$ (corner of second room). The rooms are separated by a door and agents start in the same room deterministically at positions $(1,1)$ and $(2,2)$ respectively. The door will open only when one of the switches is occupied, which means that the (Manhattan) distance between one of the agents and the switch is less than $1.5$. The full state vector contains $x,y$ locations of the two agents and binary variables to indicate if doors are open but per-agent policies are conditioned on their respective states only and the state of the door. For Sparse-Rewards Tasks, the goal was set to be deterministically at the worst case, namely $(9,9)$ and to provide a positive reward to both the agents of $100$ when reached, which means again that the (Manhattan) distance between one of the agents and the switch is less than $1.5$, a reward of 00 otherwise. The second environment, Env. (ii), was the MaMuJoCo reacher environment (Peng et al., 2021). In this environment, two agents operate the two linked joints and each space dimension is discretized over $10$ bins. Per-agent policies were conditioned on their respective joint angles only. For Sparse-Rewards Tasks, the goal was set to be randomly at the worst case, namely on position (±$(\pm 0.21,\pm 0.21)$ on the boundary of the reachable area. Reaching the goal mean to have a tip position (not observable by the agents and not discretized) at a distance less that $0.05$ and provides a positive reward to both the agents of $1$ when reached, a reward of 00 otherwise.
@@
TRPE
@@
Multi-Agent TRPO
We follow the same notation in Duan et al. (2016). Agents have independent critics $(64,64)$ Dense networks and in each epoch a dataset of $N$ trajectories is gathered for a given exploration horizon $T$ for each agent, leading to the reported number of samples. Throughout the experiment the number of epochs $e$ were set equal to $e=100$, the number of trajectories building the batch size $N=20$, the KL threshold $\delta=10^{-4}$, the maximum number of off-policy iterations set to $n_{\text{off,iter}}=20$, the discount was set to $\gamma=0.99$.
@@
Figure 4: Full Visualization of Reported Experiments.

### v3

@@ Appendix B Proofs of the Main Theoretical Results
The main empirical proof of concept was based on two environments. First, Env. (i), the so called secret room environment by [19]. In this environment, two agents operate within two rooms of a $10\times 10$ discrete grid. There is one switch in each room, one in position $(1,9)$ (corner of first room), another in position $(9,1)$ (corner of second room). The rooms are separated by a door and agents start in the same room deterministically at positions $(1,1)$ and $(2,2)$ respectively. The door will open only when one of the switches is occupied, which means that the (Manhattan) distance between one of the agents and the switch is less than $1.5$. The full state vector contains $x,y$ locations of the two agents and binary variables to indicate if doors are open but per-agent policies are conditioned on their respective states only and the state of the door. For Sparse-Rewards Tasks, the goal was set to be deterministically at the worst case, namely $(9,9)$ and to provide a positive reward to both the agents of $100$ when reached, which means again that the (Manhattan) distance between one of the agents and the switch is less than $1.5$, a reward of 00 otherwise. The second environment, Env. (ii), was the MaMuJoCo reacher environment [32]. In this environment, two agents operate the two linked joints and each space dimension is discretized over $10$ bins. Per-agent policies were conditioned on their respective joint angles only. For Sparse-Rewards Tasks, the goal was set to be randomly at the worst case, namely on position (±$(\pm 0.21,\pm 0.21)$ on the boundary of the reachable area. Reaching the goal mean to have a tip position (not observable by the agents and not discretized) at a distance less that $0.05$ and provides a positive reward to both the agents of $1$ when reached, a reward of 00 otherwise.
@@
Trust Region Pure Exploration (TRPE).
@@
Limitations of TRPE
The main limitations of the proposed methods are two. First, the Monte-Carlo estimation of single-trial objectives might be sample-inefficient in high-dimensional tasks. However, more efficient estimators of single-trial objectives remain an open question in single-agent convex RL as well, as the convex nature of the problem hinders the applicability of Bellman operators. Secondly, the plug-in estimator of the entropy is applicable to discrete spaces only, but designing scalable estimators of the entropy in continuous domains is usually a contribution per se [25].
Multi-Agent TRPO (MA-TRPO).
We follow the same notation in [6]. Agents have independent critics $(64,64)$ Dense networks and in each epoch a dataset of $N$ trajectories is gathered for a given exploration horizon $T$ for each agent, leading to the reported number of samples. Throughout the experiment the number of epochs $e$ were set equal to $e=100$, the number of trajectories building the batch size $N=20$, the KL threshold $\delta=10^{-4}$, the maximum number of off-policy iterations set to $n_{\text{off,iter}}=20$, the discount was set to $\gamma=0.99$.
@@
Figure 3: Full Visualization of Reported Experiments. Experiments with longer horizons highlight how the easier the task, the less crucial the distinction between the objectives is.